You are considering two mutual funds, Fund A and Fund B, which both return 8.75%. Fund...

  1. You are considering two mutual funds, Fund A and Fund B, which both return 8.75%. Fund A has an expense ratio of 0.70% and 12b-1 fees of 1.80%. Fund B has a front end load of 2.75%, no 12b-1 fees and an expense ratio of 0.21%. The purpose of the investment is to put your brand new baby thru college in 18 years, which is the better choice? Show your work.
Answer #1

Fund A Fund A 297.80 = 100*(1+8.75%-1.8%-0.7%)^18 425.11 =100*(1-2.75%)*(1+8.75%-0.21%)^18 Fund B Fund B should be selected.

*Please rate thumbs up